OpenStudy (wolfgirl12):

The equation x^2+y^2-4x+2y=b describes a circle. Determine the y-coordinate of the center of the circle. The radius of the circle is 7 units. What is the value of b in the equation?

OpenStudy (anonymous):

complete the squares.

OpenStudy (anonymous):

\[\left( x^2-4 x+? \right)+\left( y^2+2 y+? \right)=b+?+?\]

OpenStudy (anonymous):

\[b+?+?=7^2\] find b

OpenStudy (anonymous):

\[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] (h,k) is the center and r is the radius.

OpenStudy (anonymous):

The equation x^2+y^2-4x+2y=b describes a circle. The radius of the circle is 7 units. What is the value of b in the equation?